Please help me graph 2x^2+6x+5=0
Answer by ewatrrr(24785)

 
Hi,
Please help me graph
2x^2+6x+5=0 ||Recommend completing Square
y = 2(x+3/2)^2 - 9/2 + 5
y = 2(x+3/2)^2 +1/2 Parabola Opening Upward as 2>0, V(-3/2,1/2)
Axis of symmetry x = -3/2